Grade 12Physical Chemistry

the osmotic pressure of 3% glucose solution at 300K is ?

ketki kulkarni
8 Years agoGrade 12

1 Answer

Akash
8 Years ago
P=i CTRhere i =1C- molarity=3×1000/100=30So p=30×8.314×300P=74286here i =1C- molarity=3×1000/100=30So p=30×8.314×300P=74286here i =1C- molarity=3×1000/100=30So p=30×8.314×300P=74286